Circular No. 8953

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 L2 (HILL) R. E. Hill reports his discovery of a comet on CCD images taken with

the 0.68-m Catalina Schmidt telescope (discovery observation tabulated below), four co-added 30-s exposures on June 12 showing the new object with a 5′′–7′′ coma and a short, diffuse tail in p.a. 270^{o} in 2′′–4′′ seeing; co-added images from June 13.4 UT show a cendensed 12′′–14′′ coma with a 10′′ diffuse tail in p.a. 290^{o}, and co-added images from June 14.4 show a diffuse, fan-shaped tail ∼ 20′′ long in p.a. 250^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, J. M. Aymami (Tiana, Spain, 0.25-m reflector) writes that co-added CCD images from June 13.1 show a quasi-stellar object with a faint tail 35′′ long in p.a. 256^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
June 12.44006  23^{h}40^{m}43.^{s}93  +22^{o}42′02′′.7  18.3

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-L55.

T = 2008 July 27.949 TT  ω = 130.^{o}899
Ω = 222.914 }2000.0
q = 2.50913 AU  i = 25.341

COMETS C/2008 C6, C/2008 C7, C/2008 C8, C/2008 C9 (SOHO) Further to IAUC 8933, additional apparent comets have been found on

SOHO website images; all are Kreutz sungrazers except for C/2008 C9 (nongroup). C/2008 C6 (mag 7.5) and C/2008 C7 (mag 7) were both stellar in appearance. C/2008 C8 reached mag ∼ 5.5 and appeared teardrop-shaped in C3 images, while it was a very faint, thin streak in C2 images. C/2008 C9 was marginally diffuse and very small (mag ∼ 7.5). An asterisk in the instrumentation column below indicates that astrometry was also obtained from STEREO-A/SECCHI HI1 images (see IAUC 8926).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 C6 Feb. 7.696 21^{h}36.^{m}9 −16^{o}16′ C3* BZ 2008-L18
C/2008 C7  9.196 21 47.1 −16 17 C3* HS 2008-L18
C/2008 C8  14.821 22 17.9 −15 38 C3/2 JS 2008-L18
C/2008 C9  15.063 21 53.3 −11 33 C2  RK 2008-L18
2008 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green